In 100 meiosis, 3 crossover events occurred between genes A and B on one homologous chromosome and a, b on the other, what is the r between genes A and B
Answer #1

single Cross over event we get 2 possibilities=     A-------b and a-------B

3 cross over = 6 offspring

100-3=97 normal meiosis

Each meiosis gives 4 cell = 97 X 4 = 388

In 3 cross over event still 6 cells are in normal.

r between A and B = 6/400 = 0.015cM